OpenStudy (anonymous):

Find the exact numerical value of sin (3pi/2) + cos (pi/4)

OpenStudy (anonymous):

1.) sin(3pi/2) is -1 2.) cos(pi/4) is \[\sqrt{2}/2\] so \[(2+\sqrt{2})/2\]

OpenStudy (anonymous):

if you square the problem you get (4+2)/2 that = 3